# Continual-MEGA: A Large-scale Benchmark for Generalizable Continual Anomaly Detection

### v2

@@ Abstract
In this paper, we introduce a new benchmark for continual learning in anomaly detection, aimed at better reflecting real-world deployment scenarios. Our benchmark, Continual-MEGA, includes a large and diverse dataset that significantly expands existing evaluation settings by combining carefully curated existing datasets with our newly proposed dataset, ContinualAD. In addition to standard continual learning with expanded quantity, we propose a novel scenario that measures zero-shot generalization to unseen classes—those not observed during continual adaptation. This setting poses a new problem setting that continual adaptation also enhances zero-shot performance. We also present a unified baseline algorithm that improves robustness in few-shot detection and maintains strong generalization. Through extensive evaluations, we report three key findings: (1) existing methods show substantial room for improvement, particularly in pixel-level defect localization; (2) our proposed method consistently outperforms prior approaches; and (3) the newly introduced ContinualAD dataset enhances the performance of strong anomaly detection models. We release the benchmark and code in https://github.com/Continual-Mega/Continual-Mega.
@@ 1 Introduction
Anomaly detection (AD) cao2024adaclip ; defard2021padim ; deng2022anomaly ; gudovskiy2022cflow ; huang2024adapting ; jeong2023winclip ; liu2023simplenet ; roth2022towards ; strater2024generalad ; yao2024hierarchical ; yao2023one ; you2022unified ; zhang2023destseg ; zhou2023anomalyclip plays a critical role in quality control, ensuring precise identification of defects during production. It is widely used in sectors such as manufacturing, where automatic anomaly detection can significantly improve operational efficiency and product safety. Due to the complexity of real-world environments, anomaly detection models are required to recognize a diverse range of defects. To address the issue, many scenarios with public datasets bergmann2019mvtec ; zou2022spot ; mishra2021vt ; wang2024real ; lehr2024viaduct ; lehr2024ad3 have been proposed. Conventional deep approaches bergmann2022beyond ; cohen2020sub ; defard2021padim ; li2021cutpaste ; ristea2022self ; roth2022towards ; zavrtanik2021draem ; zou2022spot suppose unsupervised or per-class anomaly detections. Following the advancement of CLIP radford2021learning and its initial application to anomaly detection jeong2023winclip , unified anomaly detection frameworks you2022unified ; he2024learning ; yao2024hierarchical ; strater2024generalad have emerged, enabling a single model to handle diverse evaluation scenarios. These approaches include continuous learning and adaption li2022towards ; pang2025context ; liu2024unsupervised ; tang2024incremental ; jin2024oner ; meng2024moead ; mcintosh2024unsupervised as well as zero-, few-shot anomaly detection jeong2023winclip ; zhou2023anomalyclip ; li2024promptad ; zhu2024toward ; deng2023anovl ; chen2024clip ; chen2023zero ; tamura2023random ; gu2024filo ; gui2024few ; qu2024vcp .
@@
From a dataset perspective, the inherent difficulty of collecting large numbers of samples, particularly defective ones, makes anomaly detection (AD) more challenging than general vision tasks. Consequently, widely used evaluation datasets bergmann2019mvtec ; zou2022spot are notably limited in both size and variability. This limitation would have motivated recent research to explore continual, zero-shot, and few-shot learning settings as strategies to overcome the scarcity of data. Reflecting the need, we suggest the need of new evaluation benchmarks with expanded data quantity, achieved by integrating diverse public datasets and curating additional samples to increase both the volume and variety of data.
@@
In this paper, we introduce a novel and comprehensive evaluation benchmark, Continual-MEGA, that evaluates the continual and zero-shot capabilities of anomaly detection models. Our benchmark includes a large-scale evaluation dataset that integrates widely used public datasets bergmann2019mvtec ; lehr2024viaduct ; lehr2024ad3 ; mishra2021vt ; wang2024real with a newly curated dataset, ContinualAD. The Continual-MEGA dataset supports two primary evaluation scenarios: (1) a standard continual learning setup, and (2) an extended setup for evaluating generalization performance after the continual learning phase, often required in applications. To further validate the effectiveness of our curated dataset, ContinualAD, we also introduce an additional scenario where the ContinualAD dataset is excluded from training.
The extensive evaluation conducted on the proposed Continual-MEGA benchmark, we test the representative anomaly detection methods cao2024adaclip ; huang2024adapting ; liu2024unsupervised ; liu2023simplenet ; qu2024vcp ; strater2024generalad ; tang2024incremental ; tao2024kernel ; yao2024hierarchical ; zhang2024mediclip ; zhou2023anomalyclip and clearly demonstrate that there is still substantial room for improvement in the AD domain in perspective of continual adaptation and generalizability. Additionally, we propose a new baseline AD method for the Continual-MEGA benchmark that enables parameter-efficient and continuous adaptation of pre-trained CLIP. Our approach applies mixture-of-expert (MoE) style multi-layered-perceptron (MLP) adaptor utilization, combined with anomaly feature synthesis and fine-tuning of prompt embeddings, reporting the state-of-the-art performance in the proposed Continual-MEGA benchmark, expected to act as a new baseline method for future research.
Our contributions are summarized as follows:
@@
  • From extensive evaluations on the proposed Continual-MEGA benchmark, we demonstrate that there is still enough room for improvement in the performance of AD algorithms.
  • We propose a new anomaly detection method called Anomaly Detection across Continual Tasks (ADCT) that combines MoE-style adaptor modules, anomaly feature synthesis, and prompt-based feature tuning with CLIP. Our method achieves state-of-the-art performance on the Continual-MEGA benchmark and would serve as a strong baseline for future research.
@@ 2 Continual-MEGA Benchmark
We compose a new benchmark to evaluate continual learning for anomaly detection in large-scale real-world settings, including various public datasets including MVTec-AD bergmann2019mvtec , VisA zou2022spot , Real-IAD wang2024real , VIADUCT lehr2024ad3 , BTAD mishra2021vt , and MPDD jezek2021deep , and also with the newly proposed ConitnualAD dataset, which consists of diverse images collected from real-world objects. Figure 1 shows example images from the seven datasets included in the proposed Continual-MEGA benchmark, demonstrating the diversity and complexity of anomaly types across domains. The ConitnualAD dataset consists of a total of 30 classes, comprising 14,655 normal images and 15,827 anomaly images, significantly larger than widely used MVTec-AD and VisA datasets, as illustrated in Figure 2. To evaluate the continual learning performance of comparative methods, we design two experimental scenarios. The model is initially pre-trained on either 85 or 58 classes, followed by continual learning with 60 novel classes introduced incrementally.
Dataset Acquisition.
The proposed ContinualAD dataset images were collected using 10 devices: Galaxy S21+, iPhone 12 Pro Max, iPhone 13, iPhone 15 Pro Max, iPhone XS, iPad Air 4, iPad Pro 11-inch 2nd generation, iPad Pro 12.9-inch 4th generation, iPhone 12 mini, and ZFLIP 3. This diversity in devices and capturing conditions enables evaluation under a wide range of real-world anomaly scenarios and environmental variations.
@@
[…]
@@
Scenario 1 extends conventional continual learning settings by combining the MVTec-AD bergmann2019mvtec and VisA zou2022spot datasets, widely used in anomaly detection. We pretrain the model on all 858585 Base classes and sequentially introduce 555, 101010, and 303030 New classes over 121212, 666, and 222 iterations, respectively. Scenario 2 is designed to evaluate zero-shot generalization following continual adaptation. In this setting, both MVTec-AD and VisA are excluded from the continual learning process, they are neither part of the Base nor New classes. Instead, they are held out solely for assessing the model’s zero-shot performance, serving as a novel protocol to evaluate cross-domain generalization. Additionally, Scenario 3 further analyzes the generalization capability of the proposed ContinualAD dataset by removing the target dataset from the Base classes and New classes stream. Specifically, the model is continually adapted with 303030 New classes from other datasets, while zero-shot generalization is evaluated on the excluded dataset, following the setup of Scenario 2.
@@
Figure 3 demonstrates the detailed statistics for each of the two main scenarios. For both cases, we can see that there exists an imbalance for each classes. Considering that we measure the amount of forgetting class-wise, we can suppose that it would be advantageous for fitting smaller classes from previous datasets bergmann2019mvtec ; zou2022spot . We investigate the supposition by comparing the results from Scenario 1 and 2 in the Experiment section.
Figure 2: Illustration of statistics of various datasets. Each graph (from left to right) shows the number of classes, number of images, and pixel value variance for public datasets, as well as our proposed ContinualAD and Continual-MEGA.
(a) Class distribution of Scenario 1.
(b) Class distribution of Scenario 2.
Figure 3: Class distributions of different scenarios. Each colored region denotes a Base, and each divided (New) tasks. The volume in the orange line in the histogram represents the number of anomaly samples over the entire sample volume represented by gray. Detailed class configuration for each task will be presented in the Supplementary material.
Metrics and Implementation Details.
For all quantitative evaluations, we adopt two metrics proposed in tang2024incremental to evaluate continual learning performance in anomaly detection: average accuracy (ACC) and forgetting measure (FM). The ACC metrics are computed on the basis of the image-level area under the ROC curve (AUROC) and the pixel-level average precision (AP) liu2024unsupervised , providing a comprehensive view of both the classification accuracy and the model’s resilience to forgetting over time. For the FM measure, we evaluate the accuracy drop for each ACC after the adaptation. All the methods are trained and evaluated in equivalent training settings with 50 epochs for baseline method training, and 20 epochs for continual adaptation for fair comparison, following concerns raised in cha2024hyperparameters regarding the fair comparison by the hyper-parameter settings of continual learning evaluation. For the proposed baseline method, a single set of hyperparameters was derived through lightweight tuning on the base classes of Scenario 1 and then consistently applied across all remaining scenarios without further adjustment. This avoids per-scenario tuning and better reflects realistic continual learning settings, where fine-grained tuning for each incoming task is typically impractical.

### v3

@@ Abstract
We introduce a new benchmark for continual learning in anomaly detection, addressing the limited adaptability of static models in dynamic real-world deployment scenarios. Our benchmark, Continual-MEGA, expands existing evaluation settings by integrating diverse public datasets with our newly proposed dataset, ContinualAD. Beyond standard continual learning settings, we additionally propose a scenario that evaluates zero-shot generalization to unseen classes not encountered during continual adaptation. This reflects recent advances in continual zero-shot research and highlights its practical significance. This setting introduces a new agenda for the anomaly detection field, and we conduct extensive evaluations of various existing anomaly detection algorithms designed for continual or zero-shot scenarios, as well as our proposed baseline methods. From our experiments, we derive three key findings: (1) existing methods exhibit significant limitations, particularly in pixel-level defect localization, (2) the proposed ContinualAD dataset is effective for the proposed benchmarking scenario, and (3) our baseline method suggests a promising direction for designing CLIP-based continual and generalizable frameworks through simple adaptation combined with feature synthesis.
Machine Learning, ICML
@@ 1 Introduction
Anomaly detection (AD) (Roth et al., 2022; Liu et al., 2023; Jeong et al., 2023; Zhou et al., 2023) plays a crucial role in quality control, ensuring precise identification of defects during production. It is widely applied in automated detection across diverse domains, including industrial and agricultural products, medical images, and other application areas (Huang et al., 2024; Wei et al., 2025). Due to the complexity and variety of real-world environments, anomaly detection models need to recognize a wide range of defects (Guo and Lv, 2025). To tackle this issue, several benchmark scenarios have been established using public datasets (Bergmann et al., 2019; Zou et al., 2022; Mishra et al., 2021; Wang et al., 2024a; Jezek et al., 2021; Lehr et al., 2024). These datasets, such as the widely used MVTec-AD (Bergmann et al., 2019) and VisA (Zou et al., 2022), encompass a wide range of object categories, including fabrics, food items, consumer goods, and industrial components.
Conventional deep approaches (Bergmann et al., 2022; Cohen and Hoshen, 2020; Defard et al., 2021; Li et al., 2021; Ristea et al., 2022; Roth et al., 2022; Zavrtanik et al., 2021; Zou et al., 2022) assume unsupervised or per-class anomaly detection. Following the advancement of CLIP (Radford et al., 2021) and its initial application to AD (Jeong et al., 2023), unified AD frameworks (You et al., 2022; He et al., 2024; Yao et al., 2024a; Sträter et al., 2024) have emerged, allowing a single model to handle various evaluation scenarios. These approaches include continual learning and adaptation (Li et al., 2022; Pang and Li, 2025; Liu et al., 2024; Tang et al., 2024; Jin et al., 2024; Meng et al., 2024; McIntosh and Albu, 2024) as well as zero-shot, few-shot AD (Jeong et al., 2023; Zhou et al., 2023; Li et al., 2024; Zhu and Pang, 2024; Deng et al., 2023; Chen et al., 2024b, 2023; Tamura, 2023; Gu et al., 2024; Gui et al., 2024; Qu et al., 2024).
@@
From a dataset perspective, the difficulty of collecting defect samples makes AD more challenging than general vision tasks (Fang et al., 2023). Consequently, widely used evaluation datasets (Bergmann et al., 2019; Zou et al., 2022) are significantly limited in both size and variability. This limitation has motivated recent research to explore continual (Liu et al., 2024), zero-shot (Zhou et al., 2023), and few-shot learning (Zhang et al., 2024a) settings as strategies to overcome data scarcity. In light of this limitation, we suggest the need for new evaluation benchmarks with a larger quantity of data, achieved by integrating diverse public datasets and curating additional samples to increase both the volume and the variety of data.
Due to these limitations, anomaly detection systems deployed in real-world environments often face sequentially arriving tasks, where new object categories or defect types emerge over time (Li et al., 2022; Liu et al., 2024). In such scenarios, retraining models from scratch for every new task is not only computationally expensive but often infeasible due to strict data retention policies or privacy regulations that prohibit storing long-term historical data (Wang et al., 2024b; Liu et al., 2024). Furthermore, the extreme rarity of anomaly samples makes it difficult to re-collect identical defect patterns from the past, rendering catastrophic forgetting a critical failure in maintaining factory-wide safety (Bugarin et al., 2024). Continual learning aims to address these challenges by enabling the models to incrementally adapt to new data while preserving knowledge of previously seen tasks. However, in many practical cases, some tasks or defect types may not be observed during training at all, which requires models to generalize to the entirely unseen classes (i.e., tasks or defect types) (Zhou et al., 2023). This setting, known as continual zero-shot learning (CZSL) (Zhang et al., 2023), is particularly crucial to building robust and scalable anomaly detection systems, as illustrated in Figure 1 with a real-world example.
Figure 1: Motivation for using Continual Learning (CL) and Continual Zero-Shot Learning (CZSL) in anomaly detection, enabling models to handle evolving defects over time and generalize to unseen anomalies without retraining.
@@
In this paper, we introduce a novel and comprehensive evaluation benchmark, Continual-MEGA, that is designed to evaluate the continual and zero-shot capabilities of anomaly detection models. Our benchmark includes a large-scale evaluation dataset that integrates widely used public datasets (Bergmann et al., 2019; Zou et al., 2022; Mishra et al., 2021; Wang et al., 2024a; Jezek et al., 2021; Lehr et al., 2024) with a newly curated dataset, ContinualAD. The Continual-MEGA benchmark supports two primary evaluation scenarios: (1) a standard continual learning setup, and (2) an extended setup evaluating generalization performance after the continual learning phase, often required in real-world applications. This second setting aligns with the concept of continual zero-shot learning (CZSL) (Zhang et al., 2023), where models are expected to adapt to new tasks incrementally while maintaining strong generalization to entirely unseen classes. We highlight that the development of zero-shot unified AD models, an increasingly popular direction, naturally includes addressing a stream of continually incoming novel objects, which corresponds to the CZSL application scenario, as illustrated in Figure 1.
We conduct extensive evaluation on the proposed Continual-MEGA benchmark, testing representative anomaly detection methods (Cao et al., 2024; Huang et al., 2024; Liu et al., 2024, 2023; Qu et al., 2024; Sträter et al., 2024; Tang et al., 2024; Tao et al., 2024; Yao et al., 2024a; Zhang et al., 2024a; Zhou et al., 2023) and clearly demonstrating that substantial room for improvement remains for the AD domain in terms of continual adaptation and generalizability. Furthermore, through the proposed baseline method, Anomaly Detection across Continual Tasks (ADCT), aligning with the evaluation results, we show that the method employing minimal adapters with feature synthesis applied to the pretrained CLIP backbone achieves overall superior performance, suggesting that excessive adaptation and guidance to CLIP may lead to overfitting on previously seen objects. Our contributions are summarized as follows:
@@
  • From extensive evaluations on the proposed Continual-MEGA benchmark, we demonstrate that there is substantial enough room for improvement in AD performance.
  • We introduce a baseline method, Anomaly Detection across Continual Tasks (ADCT), which integrates lightweight MoE-style adapter modules and anomaly feature synthesis with CLIP, pointing toward a promising direction for both Continual and CZSL scenarios.
2 Related Works
Anomaly detection focuses on detecting and rejecting unknown samples (Amodei et al., 2016; Hendrycks et al., 2021), framed as an out-of-distribution (OOD) problem. Recent advances in large-scale backbone models, such as CLIP (Radford et al., 2021), offer a promising solution to the challenge of unified AD across categories (Jeong et al., 2023; He et al., 2024; Yao et al., 2024a; Sträter et al., 2024). Following the initial approach (Jeong et al., 2023) using CLIP, current research trends focus on developing unified anomaly detection models with zero- and few-shot category adaptation.
Zero- and Few-shot Adaptation.
Anomaly detection with zero- and few-shot adaptation (Jeong et al., 2023; Li et al., 2024; Chen et al., 2023; Tamura, 2023; Gu et al., 2024; Gui et al., 2024; Qu et al., 2024; Zhou et al., 2023; Deng et al., 2023; Chen et al., 2024b) across various categories reflects real-world scenarios where acquiring a sufficient number of samples for newly incoming categories is often infeasible, and obtaining anomaly samples is even more challenging. Various methods have been proposed to address these challenges, including text prompt utilization (Jeong et al., 2023; Li et al., 2024; Zhou et al., 2023; Deng et al., 2023; Chen et al., 2024b; Tamura, 2023; Gu et al., 2024), visual context prompting (Qu et al., 2024; Deng et al., 2023), and anomaly dataset synthesis (Chen et al., 2023, 2024a). Notably, most of these approaches leverage text prompt information, with strategies ranging from manually designed templates (Jeong et al., 2023; Deng et al., 2023), normal-sample-only strategies (Li et al., 2024), learned prompts (Zhou et al., 2023; Deng et al., 2023; Gu et al., 2024), to augmented prompting techniques (Tamura, 2023).
Continual Adaptation.
Another notable trend in recent anomaly detection research is continual adaptation (Li et al., 2022; Pang and Li, 2025; Liu et al., 2024; Tang et al., 2024; Jin et al., 2024; Meng et al., 2024; McIntosh and Albu, 2024), where object categories arrive incrementally. In this context, the primary goal is to mitigate catastrophic forgetting while ensuring that adaptation to previous categories improves the model’s performance for future category adaptations. Based on initial efforts (Li et al., 2022), many approaches have been proposed, including context-aware feature adaptation (Pang and Li, 2025), learned text prompts (Liu et al., 2024), unified reconstruction-based detection frameworks (Tang et al., 2024), online replay memory mechanisms (Jin et al., 2024), parameter-efficient tuning strategies (Meng et al., 2024), and unsupervised tuning approaches (McIntosh and Albu, 2024; Tang et al., 2024). The continual evaluation scenario is built on public datasets such as MVTec-AD (Bergmann et al., 2019) or VisA (Zou et al., 2022), but the quantity and diversity of the dataset are limited compared to the scenario (Bang et al., 2021) using ImageNet (Deng et al., 2009).
Continual Zero-shot Learning.
Continual zero-shot learning (CZSL) (Zhang et al., 2023; Chaudhry et al., 2018) has recently emerged as a paradigm that aims to simultaneously preserve past knowledge and adapt to future tasks, mirroring the way humans learn throughout their lifetime. Subsequent studies have addressed these problems by generative replay (Gautam et al., 2024), context composition (Zhang et al., 2024b), and class normalization (Skorokhodov and Elhoseiny, 2020). Recognizing the practical relevance of the CZSL setting in AD applications, we construct a large-scale Continual-MEGA benchmark designed to incorporate CZSL scenarios.
Figure 2: Illustration of statistics of various datasets. Each graph (from left to right) shows the number of classes, number of images, and pixel value variance for public datasets, as well as our proposed ContinualAD and Continual-MEGA. Image variance is defined as the average per-pixel variance for each class.
@@ 3 Continual-Mega Benchmark
3.1 ContinualAD Dataset
To form the Continual-MEGA benchmark, we propose the ContinualAD dataset, which is significantly larger in scale compared to previous datasets. The ContinualAD dataset comprises 30 classes; a detailed breakdown of the normal and anomaly sample counts for each category is summarized in Table A in the Appendix. Moreover, the ContinualAD dataset includes a wide range of object instances within the same class, enabling the evaluation of robustness to intra-class variation. Consequently, as illustrated in Figure 2, the ContinualAD dataset exhibits significantly higher image variance than existing datasets. The variance value is computed by first calculating the pixel-wise variance across images within each class and then averaging these values across all classes in the dataset. This indicates that prior datasets primarily contain visually similar images within each class, limiting their ability to evaluate model performance under diverse conditions. In contrast, the higher variance in ContinualAD facilitates more realistic and challenging evaluation scenarios.
Dataset Acquisition.
The proposed ContinualAD dataset is constructed from 30 real-world object categories, for which normal images were first collected under normal conditions. For each object, we then deliberately induced diverse defects such as cracks, holes, rot, scratches, bending, and contamination to obtain corresponding anomaly images. All images were captured using 10 devices (Galaxy S21+, iPhone 12 Pro Max, iPhone 13, iPhone 15 Pro Max, iPhone XS, iPad Air 4, iPad Pro 11-inch 2nd generation, iPad Pro 12.9-inch 4th generation, iPhone 12 mini, and ZFLIP 3), and the anomaly regions in the anomaly images were annotated at the pixel level using polygon masks. This diversity in devices and capturing conditions enables evaluation under a wide range of real-world anomaly scenarios and environmental variations.
Dataset Statistics.
Figure 2 further summarizes the overall class distribution and the number of samples per class. The graph highlights that the ContinualAD dataset offers a competitively large number of classes and samples compared to existing Real-IAD dataset. Notably, the scope of the proposed Continual-MEGA benchmark extends beyond IAD, including a diverse range of anomalies and object types, similar to the variety found in widely used datasets such as MVTec-AD (Bergmann et al., 2019). The ContinualAD dataset consists of a total of 30 classes, comprising 14,655 normal images and 15,827 anomaly images, notably larger than widely used MVTec-AD and VisA datasets.
3.2 Benchmark Configuration
@@
We compose a new benchmark to evaluate continual learning for anomaly detection in large-scale real-world settings, comprising various public datasets including MVTec-AD (Bergmann et al., 2019), VisA (Zou et al., 2022), Real-IAD (Wang et al., 2024a), VIADUCT (Lehr et al., 2024), BTAD (Mishra et al., 2021), and MPDD (Jezek et al., 2021), and the newly proposed ContinualAD dataset, which consists of diverse images collected from real-world objects. Figure 3 shows example images from the seven datasets included in the proposed Continual-MEGA benchmark. These examples highlight the diversity and complexity of anomaly types across domains. To evaluate the continual learning performance, we design two experimental scenarios. The model is initially pre-trained on either 85 or 58 classes, followed by continual learning with 60 novel classes introduced incrementally.
Figure 4: Overview architecture of training process. Synthetic feature generation is performed exclusively during the training phase. At inference, adapters trained for each task are accumulated for evaluation. A detailed overview of the inference stage is provided in Appendix Figure B.
@@
To construct a large-scale continual learning benchmark for anomaly detection, we integrate seven datasets into three distinct evaluation scenarios. In each scenario, the continual learning setup is denoted as $(\#\text{Base})$-$(\#\text{New})$, where $(\#\text{Base})$ and $(\#\text{New})$ represent the number of base and newly introduced classes, referred to as Base and New, respectively. The first two scenarios, Scenario 1 and Scenario 2, represent the main evaluation of the benchmark. Furthermore, to evaluate the effectiveness of the proposed ContinualAD dataset, we conduct Scenario 3 to compare it with Scenario 2.
@@
Scenario 1 extends conventional continual learning settings by combining the MVTec-AD (Bergmann et al., 2019) and VisA (Zou et al., 2022) datasets, widely used in anomaly detection. We pretrain the model on all $85$ Base classes and sequentially introduce $5$, $10$, and $30$ New classes over $12$, 66, and $2$ iterations, respectively. Scenario 2 is designed to evaluate zero-shot generalization following continual adaptation. In this setting, both MVTec-AD and VisA are excluded from the continual learning process, as they are neither part of the Base nor New classes. Instead, they are held out solely for assessing the model’s zero-shot performance, serving as a novel protocol to evaluate cross-domain generalization. Additionally, Scenario 3 further analyzes the generalization capability of the proposed ContinualAD dataset by removing the target dataset from the Base classes and New classes stream. Specifically, the model is continually adapted with $30$ New classes from other datasets, while zero-shot generalization is evaluated on the excluded dataset, following the setup of Scenario 2. An overview of the three scenarios, including the base and new classes and the held-out datasets, is provided in Table B of the Appendix.
@@
Detailed class distributions for Scenarios 1, 2, and 3 are illustrated in Appendix Figures DE, and F, respectively. These figures reveal a significant class imbalance across the datasets. Considering that we measure the amount of forgetting class-wise, we hypothesize that better fitting smaller classes from previous datasets (Bergmann et al., 2019; Zou et al., 2022) could be advantageous. We further investigate this supposition by comparing the results across the three scenarios in the Experiments section.
Metrics.
For all quantitative evaluations, we adopt two metrics proposed in (Tang et al., 2024) for evaluating continual learning performance in anomaly detection: average accuracy (ACC) and forgetting measure (FM). The ACC metrics are computed on the basis of the image-level area under the ROC curve (AUROC) and the pixel-level average precision (AP) (Liu et al., 2024), providing a comprehensive view of both the classification accuracy and the model’s resilience to forgetting over time. For the FM measure, we measure the decrease in ACC after adaptation. To characterize the overall performance in more detail, we additionally report the mean ACC and FM over the image- and pixel-level scores, providing a compact summary of the joint performance across both granularities.
Training Protocol and Fair Comparison.
All methods, including our proposed baseline, are trained under the same continual learning protocol in every scenario. Within each scenario, we use the same set of base classes and the same task order across all methods. For each method, the base classes are trained for 50 epochs and each continual adaptation phase is trained for 20 epochs, ensuring identical training budgets. Following the concerns in (Cha and Cho, 2024) about fair comparison in continual learning, we avoid per-scenario hyper-parameter tuning: for existing baselines, we use the publicly available default hyper-parameters and data augmentation settings from the authors’ implementations or papers, while for our proposed baseline method a single set of hyper-parameters is obtained via lightweight tuning on the base classes of Scenario 1 and then fixed for all remaining scenarios. Moreover, to make the comparison more conservative, we do not apply any additional data augmentation for our baseline beyond the default preprocessing pipeline, whereas the compared methods use their standard data augmentation settings. This setup enables a controlled and fair assessment of continual learning performance across all methods.